OpenStudy (anonymous):

Jolie can read around 14 sentences per minute of her current novel. She wants to finish her reading before the movie comes out next week. If she has 17 pages left to read and each page has about 120 sentences, how many hours should Jolie set aside to complete her reading? Please show full work.

OpenStudy (anonymous):

2 hours

OpenStudy (anonymous):

17 pages x 120 sentences = 2,040 sentences 2,040/14 = minutes needed About 146 minutes needed 146 ÷ 60 = hours and minutes left over. 2 hours and 26 minutes is the number of hours she should set aside.

OpenStudy (anonymous):

1 page= 120 sentence so divide the total of sentences of how much she can read, 120/14=8.6 and now we multiple the answer with the pages, 8.6 times the number of pages (17) and you will get, 8.6 x 17 = 146.2 which can be 2 hours and 26.2 minutes.